Over the last 3 decades, pieces have come together:

• Hadrons are composite objects made up of quarks

• 3 “generations” of quarks and leptons

( \begin{matrix}u \\ d \end{matrix} ) ( \begin{matrix}c \\ s \end{matrix} ) ( \begin{matrix}t \\ b \end{matrix} )
( \begin{matrix}e \\ νe \end{matrix} ) ( \begin{matrix}μ \\ νμ \end{matrix} ) ( \begin{matrix}τ \\ ντ \end{matrix} )

⇔ ( \begin{matrix}Iz = +1/2 \\ Iz = −1/2 \end{matrix} )

• Three of the four forces can be described by relativistic quantum field theories:

– “bosons” mediate each force

EM  γ
Weak  W+ , W− , Z∘
Strong gluons